OpenStudy (anonymous):

how would i solve (3cos(3theta))^2

OpenStudy (anonymous):

by solve i mean distribute

OpenStudy (anonymous):

i know 3^2=9 but not sure how to handle cos(3theta) what do i do with the 3theta?

OpenStudy (jdoe0001):

\(\bf [3cos(3\theta)]^2\implies 3^2[cos(3\theta)]^2\implies 9cos^2(3\theta) \\ \quad \\ \textit{triple angle identity}\to{\color{brown}{ cos^3(\theta)= 4cos^3(\theta)-3cos(\theta)}} \\ \quad \\ 9cos^2(3\theta)\implies 9[4cos^3(\theta)-3cos(\theta)]^2\) is what I can see thus far if you're just distributing, well.. you can just expand that like you would any perfect square trinomial

OpenStudy (anonymous):

@jdoe0001 it is actually part of a a problem im having. finding area of a polar equation. r=3cos(3\Theta)

OpenStudy (anonymous):

supposed to be r=3cos(3theta)